# EFSA: Towards Event-Level Financial Sentiment Analysis

### v1

@@ 5.1 Main Results and Analysis
General Domain LLMs. LLMs (ChatGPT, GPT-4, and GLM-4) under zero-shot settings fail to produce satisfactory results. Few-shot settings can enhance their performance, with particularly noticeable improvements for event classification tasks and slight improvements for sentiment analysis tasks. This may be attributed to the inherent capabilities of larger-parameter LLMs in sentiment analysis. Few-shot demonstrations primarily boost the LLMs’ proficiency in our specialized event classification tasks.
Fine-tuning LLMs can significantly improve performance scores by enhancing both domain-specific ability and capability for structured output. Smaller-parameter LLMs (e.g. ChatGLM3-6B) under zero-shot settings tend to produce outputs at will, often deviating from the instructions of structured output, which presents substantial challenges in statistically scoring model performance and poor performance scores.
@@
Financial Domain-Specific LLMs. Financial domain-specific LLMs are fine-tuned or pre-trained on open-source financial domain datasets based on general domain LLMs. Our benchmark results exhibit the enhanced performance of domain-specific LLMs relative to the original base models under the zero-shot settings (DISC vs. Baichuan, Xuanyuan vs. Llama), suggesting that domain-specific can significantly enhance their domain-specific ability. Furthermore, domain-specific LLMs demonstrate a commendable capability for structured output, which may be attributed to other structured tasks within their training datasets. However, financial domain-specific LLMs do not outperform LLMs that are fine-tuned on our dataset. Our dataset can serve as a rich resource to facilitate the advancement of financial domain-specific LLMs.
@@
SLMs. The benchmark scores of SLMs underscore that SLMs outstrip the performance of zero-shot LLMs, affirming the conclusions presented in Zhang et al. (2023)’s study. However, due to the constraints imposed by task difficulty, the quantity of parameters has delimited the upper potential of the model. Consequently, the performance of SLMs does not surpass fine-tuned LLMs.
@@
Our Framework. Our Chain of Thought (CoT) framework augments the efficacy of non-open-source LLMs through prompt-based adjustments alone. For open-source LLMs, the dialogue fine-tuning empowers the foundational model to realize more substantial score enhancements, particularly notable in models with initially subpar performance (e.g., Llama). Notably, the CoT framework hurts entity-level FSA tasks. This negative influence may be because the comprehensive CoT sequence spans from company identification to event prediction and ultimately to sentiment analysis, while direct inference from company to sentiment may lead LLMs astray, culminating in counterproductive effects.
@@ 5.2 Case Study
Figure 4: Examples provided include the input news text, the corresponding gold labels, and the predicted quantities. The red font denotes the incorrect part of the prediction
To better demonstrate the effectiveness of our framework, we perform a case study on GPT-4, as shown in Figure 3. As shown in Example 1, GPT-4 correctly predicted fine-grained events but made mistakes in predicting coarse-grained events, confusing the two difficult-to-distinguish categories of shareholder affairs and stock affairs. It is possibly due to the multiple occurrences of stock in the text leading to misdirection. Under our CoT framework, in the second hop of reasoning from company to coarse-grained events, GPT-4 focuses more on events directly related to the company itself, thereby making correct judgments. The CoT framework can prevent a situation where the predictions for fine-grained events are correct, yet those for coarse-grained events are not. For sentiment prediction, an inexperienced LLM can easily be misled by the company receives a notice of share reduction. However, under the CoT framework, GPT-4 pays more attention to the sentiment of the fine-grained event (Stock Holding Adjustment) itself and can deduce the correct sentiment from the subsequent information that there was no actual share reduction action.
Error Analysis. We observed that LLMs sometimes produce outputs beyond the defined label sets, specifically illustrated in example-2 and -3 from our error study. LLMs may alter the fixed label instead of outputting as instructed. This observation aligns with Zhang et al. (2021a)’s previous research, highlighting the nature of the generation modeling since it does not perform “extraction” in the given sentence. This phenomenon is more pronounced in LLMs with smaller parameter sizes and can be mitigated through fine-tuning. Similar to instructing LLMs toward structured outputs, fine-tuning significantly enhances the ability of smaller-parameter LLMs to output as required.
@@ 6 Related Work
Previous research on the FSA dataset has concentrated on sentence or document level (Takala et al., 2014; Malo et al., 2014; Cortis et al., 2017; Sinha et al., 2022). This is based on an assumption that the given text conveys a single sentiment towards a certain topic. Recent FSA datasets exhibit a trend towards progressively finer granularity. However, fine-grained datasets (Maia et al., 2018; Tang et al., 2023) mostly focus on entities and sentiments, neglecting the concern of events. Furthermore, the resources for fine-grained FSA datasets are still limited (Du et al., 2024). Following Du et al. (2024), we summarize the most widely used and recent FSA benchmark datasets in Appendix C
@@ 7 Conclusion
In this paper, we present EFSA, a novel task for Financial Sentiment Analysis (FSA), which deepens FSA to the event level. To support this task, we constructed the largest annotated Chinese dataset from a large-scale financial news corpus. We evaluated this task on widely used both small and large language models to present benchmark scores on the proposed dataset. Additionally, we designed a 4-hop reasoning prompting framework based on existing LLMs to resolve this task. Our experiments demonstrate the challenge of EFSA and the effectiveness of our proposed approach. Our work opens a new avenue in FSA and offers significant value to the entire financial domain.
@@
Due to budget constraints, we only conducted experiments on a subset of our dataset using LLMs that have not been open-sourced yet(Chatgpt, GPT-4, GLM-4), on a single run. This may result in bias in our evaluation scores.
@@ Appendix A Event Taxonomy
The event taxonomy is translated from Chinese; for more accurate and detailed information and the full text of annotation guidance, please refer to our open-source website. https://anonymous.4open.science/r/EFSA-645E
@@
Appendix B Prompt
@@
Appendix C FSA Benchmark Datasets
Table 3 summarize the most widely used and recent FSA benchmark datasets

### v2

@@ 5.1 Main Results and Analysis
General Domain LLMs. Our prior experiments demonstrate that smaller-parameter LLMs (ChatGLM, Llama2-Chinese, Baichuan2, and QwenLM) perform poorly in zero-shot and few-shot settings (c.f. Appendix D). This can be attributed to the inherent complexity of the EFSA task and the limited capacity of smaller models to produce structured outputs, which leads to their failure in generating the specified quadruple responce format, culminating in incorrect outputs. Fine-tuning these LLMs can significantly improve their performance scores by enhancing both domain-specific ability and capability for structured output.
@@
Larger-parameter LLMs (ChatGPT, GPT-4, and GLM-4) under zero-shot settings also fail to produce satisfactory results. Few-shot settings can enhance their performance, with particularly noticeable improvements for event classification tasks and slight improvements for sentiment analysis tasks. This may be attributed to the inherent capabilities of larger-parameter LLMs in sentiment analysis tasks. Few-shot demonstrations primarily boost the LLMs’ proficiency in our specialized event classification tasks.
@@
Financial Domain-Specific LLMs. Financial domain-specific LLMs are fine-tuned or pre-trained on open-source financial domain datasets based on general domain LLMs. Our benchmark results exhibit the enhanced performance of domain-specific LLMs relative to their original base models under the zero-shot settings (DISC vs. Baichuan, Xuanyuan vs. Llama), suggesting that domain-specific customization can significantly enhance performance on previously unseen tasks within the same domain. Furthermore, domain-specific LLMs demonstrate a commendable capability for structured output, which may be attributed to other structured tasks within their training datasets.
However, financial domain-specific LLMs do not outperform LLMs that are fine-tuned on our dataset. So we further fine-tuned financial domain-specific LLMs on our dataset. The DISC model itself is based on LoRA fine-tuning. We utilized a blend ratio of 7:3, integrating the fine-tuning weights derived from our dataset with DISC’s original weights. LoRA fine-tuning is not applicable to XuanYuan, as XuanYuan is re-pretrained. The results show that further fine-tuning can enhance the overall capabilities. Yet, it still cannot surpass the performance of their base models fine-tuned solely on our dataset. Our dataset can serve as a rich resource to facilitate the advancement of financial domain-specific LLMs.
@@
SLMs. The benchmark scores of SLMs underscore that SLMs outstrip the performance of zero-shot LLMs, affirming the conclusions presented in Zhang et al. (2023)’s study: while LLMs have shown proficiency in many sentiment analysis tasks, they fall short when it comes to extracting structured sentiment and opinion information. Smaller models have learned better structured output capabilities through fine-tuning. However, due to the constraints imposed by task difficulty, the number of parameters has limited the upper potential. Consequently, the performance of SLMs does not exceed that of fine-tuned LLMs.
@@
Our Framework. Our Chain of Thought (CoT) framework augments the efficacy of non-open-source LLMs through prompt-based adjustments alone. For open-source LLMs, the dialogue fine-tuning empowers the foundational model to realize more substantial score enhancements, particularly notable in models with initially subpar performance (e.g., Llama). Notably, the CoT framework hurts entity-level FSA tasks. This is because the score of the (company, sentiment) tuple is directly calculated based on the tuple part of the quadruple outputted by the 4-hop CoT, the process is susceptible to error accumulation. Inaccuracies in event prediction can lead to erroneous sentiment predictions, adversely affecting the accuracy of Entity-Level FSA.
@@ 5.2 Case Study
Figure 4: Examples provided include the input news text, the corresponding gold labels, and the predicted quantities. The red font denotes the incorrect part of the prediction.
To better demonstrate the effectiveness of our framework, we perform a case study on GPT-4, as shown in Figure 4. As shown in Example 1, GPT-4 correctly predicted fine-grained events but made mistakes in predicting coarse-grained events, confusing the two difficult-to-distinguish categories of shareholder affairs and stock affairs. It is possibly due to the multiple occurrences of stock in the text leading to misdirection. Under our CoT framework, in the second hop of reasoning from company to coarse-grained events, GPT-4 focuses more on events directly related to the company itself, thereby making the correct prediction. The CoT framework can prevent a situation where the predictions for fine-grained events are correct, yet those for coarse-grained events are not. For sentiment prediction, an inexperienced LLM can easily be misled by the company receives a notice of share reduction. However, under the CoT framework, GPT-4 pays more attention to the sentiment of the fine-grained event (Stock Holding Adjustment) itself and can deduce the correct sentiment from the subsequent information that there was no actual share reduction action.
Error Analysis. We observed that LLMs sometimes produce outputs beyond the defined label sets, specifically illustrated in Example-2 and -3. LLMs may alter the fixed label instead of outputting as instructed. This observation aligns with Zhang et al. (2021a)’s previous research, highlighting the nature of the generation modeling since it does not perform “extraction” in the given sentence. This phenomenon is more pronounced in LLMs with smaller parameter sizes and can be mitigated through fine-tuning. Similar to instructing LLMs toward structured outputs, fine-tuning significantly enhances the ability of smaller-parameter LLMs to output as required.
@@ 6 Related Work
Previous research on the FSA datasets has concentrated on sentence or document level (Takala et al., 2014; Malo et al., 2014; Cortis et al., 2017; Sinha et al., 2022). This is based on an assumption that the given text conveys a single sentiment towards a certain topic. Recent FSA datasets exhibit a trend towards progressively finer granularity. However, fine-grained datasets (Maia et al., 2018; Tang et al., 2023) mostly focus on entities and sentiments, neglecting the concern of events. Furthermore, the resources for fine-grained FSA datasets are still limited (Du et al., 2024). Following Du et al. (2024), we summarize the most widely used and recent FSA benchmark datasets in Appendix E.
Event-level sentiment analysis is designed to identify user emotions on social platforms regarding current events (Zhang et al., 2022a). In this paper, we broaden the scope of event-level sentiment analysis by applying it to FSA, enhancing its relevance and utility in financial contexts.
@@ 7 Conclusion
In this paper, we present EFSA, a novel task for Financial Sentiment Analysis (FSA), which deepens FSA to the event level. To support this task, we constructed the largest Chinese dataset annotated for event-level FSA from a large-scale financial news corpus. We evaluated this task on widely used language models to present benchmark scores on the proposed dataset. Additionally, we designed a 4-hop reasoning prompting framework based on existing LLMs to resolve this task. Our experiments demonstrate the challenge of EFSA and the effectiveness of our proposed approach. Our work opens a new avenue in FSA and offers significant value to the entire financial domain.
@@
Due to budget constraints, we only conducted experiments on a subset of our dataset using LLMs that have not been open-sourced yet (Chatgpt, GPT-4, GLM-4), on a single run. This may result in bias in our evaluation scores. Although CoT’s gradual processing approach leads to more accurate and reliable results, this may increase some additional computation cost and time in solving complex FSA problems.
@@
Acknowledgements
The research work is supported by the National Key R&D Plan No. 2022YFC3303305. Xiang Ao is also supported by the Project of Youth Innovation Promotion Association CAS, and Beijing Nova Program 20230484430.
@@ Appendix A Event Taxonomy
The event taxonomy is translated from Chinese; for more accurate and detailed information, please refer to our open-source website.
@@
Appendix B Annotation Platform
Figure 5 presents a screenshot of the annotation system, illustrating its primary functions.
Appendix C Prompt
The prompts are translated from Chinese; for more accurate and detailed information please refer to our open-source website.
@@ Appendix C Prompt
Appendix D Detailed Experimental Results
We conducted evaluations of all LLMs under zero-shot and three-shot settings. The performance of smaller-parameter LLMs counts was found to be suboptimal. Consequently, we excluded these results from the primary experimental results Table 2. These detailed results are reported in Table 3.
Appendix E FSA Benchmark Datasets
Table 4 summarizes the most widely used and recent FSA benchmark datasets.